Andrea saved Php370 from her allowance with denominations of 20-peso and 50-peso bills. She had 14 pieces of bills in all. How many did bills of each denomination she have?
Answer by ikleyn(52792)

.

Let x be the number of the 50-peso bills;

then the number of the  20-peso bills is  (14-x), OBVIOUSLY.



Then you write the total money equations

    50x + 20*(14-x) = 370  peso.



Next, you simplify and solve this basic equation

    50x + 280 - 20x = 370

    50x - 20x = 370 - 280

       30x    =     90

         x    =     90/30 = 3.



ANSWER.  3 50-peso bills and the rest, 14-3 = 11  20-peso bills.


CHECK.   3*50 + 11*20 = 150 + 220 = 370  peso, total.   ! Correct !

The problem is just solved.

---------------

To see millions of coin problems,  solved and carefully explained,  look into the lessons
    - Coin problems
    - More Coin problems
    - Solving coin problems without using equations
    - Kevin and Randy Muise have a jar containing coins
    - Typical coin problems from the archive
    - Three methods for solving standard (typical) coin word problems
    - More complicated coin problems
    - Advanced coin problems
    - Solving coin problems mentally by grouping without using equations
    - Non-typical coin problems
    - Santa Claus helps solving coin problem
    - OVERVIEW of lessons on coin word problems
in this site.

You will find there the lessons for all levels - from introductory to advanced,
and for all methods used - from one equation to two equations and even without equations.

A convenient place to quickly observe these lessons from a  "bird flight height"  (a top view)  is the last lesson in the list.

Read them and become an expert in solution of coin problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Coin problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.